In each parts (a)-(c), we have given a likely range for the observed value of a sample proportion p. Based on the given range, identify the educated guess that should be used for the observed value of p to calculate the required sample size for a prescribed confidence level and margin of error.
a. 0.4 to 0.7
b. 0.7 or greater
c. 0.7 or less
d. In each of parts (a)-(c), which observed values of the sample proportion will yield a largin margin of error than the one specified if the educated guess is used for the sample size computation?

Answer by stanbon(75887)

In each parts (a)-(c), we have given a likely range for the observed value of a sample proportion p.
Based on the given range, identify the educated guess that should be used for the observed value of p to calculate the required sample size for a prescribed confidence level and margin of error.
a. 0.4 to 0.7
Edu. Guess: p = (0.7+0.4)/2 = 0.55
----------------------------------
b. 0.7 or greater
Edu. Guess: (1+0.7)/2 = 0.85
--------------------------
c. 0.7 or less
Edu. Guess: (0+0.7)/2 = 0.35
==========================================
d. In each of parts (a)-(c), which observed values of the sample proportion will yield a larger margin of error than the one specified if the educated guess is used for the sample size computation?
????
============
Cheers,
Stan H.